An excess of dilute sulphuric acid reacts with both aqueous barium hydroxide and aqueous barium chloride. In what way are the two reactions the same?
1
Expert's answer
2020-05-07T14:15:33-0400

Reaction with Barium Chloride

H2SO4 + BaCl2 = BaSO4 + 2HCl


Reaction with Barium Hydroxide

Ba(OH)2 + H2SO4 = BaSO4 + 2H2O


The similarity is that in both the cases H2SO4 behaves as an strong acid.